OpenStudy (anonymous):

27^(2x)=9^(x-3), solve for x

OpenStudy (anonymous):

where x is a real number

OpenStudy (lgbasallote):

if i told you that 27 = 3^3 and 9 = 3^2 would that help?

OpenStudy (anonymous):

I just want to know what method to use to solve this. You know I'm no good at math...

OpenStudy (lgbasallote):

you use the property \[\huge a^x = a^y \implies x = y\]

OpenStudy (lgbasallote):

is that what you're asking?`

OpenStudy (lgbasallote):

@haha31896 tell me if you have an answer and want it checked

OpenStudy (anonymous):

\[ \huge 27^{(2x)}=9^{(x-3)} \implies 3^ {6x} = 3^{2x-6} \] \[\huge \implies 6x = 2x-6\implies x = -\frac 3 2\]

OpenStudy (anonymous):

Okay. I understand..thanks @lgbasallote! That formula you gave helped!

OpenStudy (lgbasallote):

welcome ^_^